# Conghui He4, Xiaoyang Wang6, Fan Wu1

### v1

@@ Abstract
Large language model (LLM) and agent techniques for data analysis (a.k.a LLM/Agent-as-Data-Analyst) have demonstrated substantial impact in both academica and industry. In comparison with traditional rule or small-model based approaches, (agentic) LLMs enable complex data understanding, natural language interfaces, semantic analysis functions, and autonomous pipeline orchestration. The technical evolution further distills five key design goals for intelligent data analysis agents, namely semantic-aware design, modality-hybrid integration, autonomous pipelines, tool-augmented workflows, and support for open-world tasks. From a modality perspective, we review LLM-based techniques for $(i)$ structured data (e.g., table question answering for relational data and NL2GQL for graph data), $(ii)$ semi-structured data (e.g., markup languages understanding and semi-structured table modeling), $(iii)$ unstructured data (e.g., chart understanding, document understanding, programming languages vulnerable detection), and $(iv)$ heterogeneous data (e.g., data retrieval and modality alignment for data lakes). Finally, we outline the remaining challenges and propose several insights and practical directions for advancing LLM/Agent-powered data analysis.
@@
Figure 2: Technical Overview of LLM/Agent-as-Data-Analyst.
@@ 1.2 Opportunities by LLM/Agent-As-Data-Analyst
O4: Autonomous Evolving. LLM agents function not as static tools but as adaptive systems capable of refining their performance through continuous interaction and learning. Such adaptability is essential, as analytical requirements and data environments inevitably evolve over time hong2024datainterpreter . For instance, a financial analysis agent may initially concentrate on trend forecasting but, through iterative feedback, expand its capabilities to detect anomalies in real time or incorporate emerging risk factors. This capacity for autonomous evolution ensures that the system remains advanced, alleviates the need for manual reconfiguration, and delivers increasingly personalized and efficient analytical support atreja2025alasautonomouslearningagent .
@@ 1.3 Techniques of LLM-Powered Data Analysis
As shown in Figure 1 and Figure 2, given the diversity of data formats and application scenarios, we establish a taxonomy of LLM-powered data analysis which progresses along two dimensions: (1) the range of data modalities supported, i.e., structured, semi-structured, unstructured, and heterogeneous (the x-axis), and (2) the evolution of interaction paradigms, i.e., code-based, domain-specific language (DSL)-based, and natural language (NL)-based (the y-axis).
@@
LLM/Agent for Structured Data Analysis. Structured data, such as relational databases relationaldata and graphs graphdata , remains central to industrial practice due to its standardized schema and well-defined semantics. Traditional approaches rely on SQL or similar code-based query languages, later extended by DSLs that simplify domain-specific tasks neo4j . With LLMs, users to interact with structured data through NL-interface, further through code generation, DSL mapping, or LLM-based question answering sqlsurvey ; shi2024sqlsurvey , with agents further orchestrating multi-step analytical workflows zhu2024sqlsurvey . The core of LLM driven structured data analysis is to use LLM for pipeline or end-to-end processing.
@@
$\bullet$ Relational Data. Relational data is typically stored and managed through specialized databases. A method for LLM-based relational data analysis is to leverage the LLM as the interface to convert analysis task in natural language into SQL or code. Then manipulate the database through SQL or code to get the analysis results. To better align natural and specialized languages, techniques like schema linking zhang2024finsql , information retrieval DINSQL and task decomposition hong2024datainterpreter are applied to the pipeline; or the model is fine-tuned yin2022code on certain tasks for better end-to-end generation. Another direction is to skip the database manipulation and directly use LLM for semantic analysis. This requires the model to better understand intentions of analysis tasks, and relationships within structured data. Hence RAG patnaik2024cabinet , prompt engineering tatllm and task decomposition zhao2024tapera are used in pipelines for semantic analysis, and MLLM zheng2024multimodal or LLM su2024tablegpt2 are trained or fine-tuned for reasoning based on textual or visual table representations. Time series data, as a specialized form of relational data characterized by temporal dependencies, can also benefit from similar LLM-powered techniques. Recent works alnegheimish2024can ; DBLP:conf/iclr/0005WMCZSCLLPW24 adopt analogous strategies, such as translating natural language into code, retrieving and transforming sequences, or employing LLMs for direct temporal reasoning.
@@
$\bullet$ Graph Data. Graph data represents entities and inter-dependencies to model complex network semantics, presenting challenges due to vast search space and complex path reasoning. There are databases and specified query language for graph data  blazegraph ; graphdb . So LLM can also serve as the interface between natural language and graph query language. Techniques like agent liang2024natnl2gql , fine-tuning and prompt engineering liang2024aligning are introduced for LLMs’ better comprehension of the graph structure and language syntax. And for semantic analysis based on the graph, The mainstream methods are reflect the ideas of RAGzhang2022subgraph , agent luo2025kbqa and fine-tuning ye2024instructglm .
@@
$\bullet$ Semi-Structured Tables represent a flexible form of tabular data whose evolution is closely tied to advances in structured table understanding and the semantic capabilities of LLMs. Traditional approaches, often based on PLMs(herzig2020tapas RoBERTa ), struggle to capture the complex layouts, irregular headers, and hierarchical structures due to their limited ability to encode semantics. With the advent of LLMs, new paradigms have emerged for handling such tables, including SemiStru2Stru transformations that convert semi-structured tables into relational-like forms 10.1145/3726302.3730071 ; li2023autotablessynthesizingmultisteptransformations , and DSL-based modeling that leverages domain-specific languages to explicitly encode structure and operations tang2025st .
LLM/Agent for Unstructured Data Analysis. Unstructured data encompass a wide range of modalities such as chart, video, document and 3D models, which lack a fixed schema and therefore pose significant challenges for conventional analysis pipelines.
@@
$\bullet$ Video. Video encodes spatial content that evolves over time, requiring joint modeling of semantics and dynamics. Conventional methods rely on vision backbones with temporal pooling or attention, which struggle with annotation costs and long-sequence efficiency chen2024timemarker wang2024grounded . Recent advances reformulate videos as structured token sequences for LLM reasoning. Agents further decompose queries into temporal grounding, multimodal fusion, and summarization, enabling richer temporal understanding and efficient computation chen2024timemarker deng2025seq2time . Building upon this, LLMs have also been extended to Video Emotional Analysis , where multimodal fusion of visual, acoustic, and textual cues enables affective state inference.In addition, posture-based emotion analysis and 3D mesh reconstruction support nuanced social interaction modeling and relational emotional predictions.jadhav2023ai ; muller2024predicting . For Object Detection, integrating high-precision detectors with multimodal reasoning allows object-centric summarization and reference grounding in videos to make correct object detections in videos de2024video ; yuan2025videorefer . Furthermore, Gesture and Behavior Detection leverages LLM-driven pipelines extract fine-grained motion features and gestures, supporting interaction analysis and embodied behavior reasoning, often via 3D reconstructions for higher fidelity whitehead2025utilizing ; weng2025artificial .
$\bullet$ Document. Documents, such as PDFs, web pages, and scanned reports, serve as primary carriers of information in business and academia. Traditional document analysis relies on Optical Character Recognition (OCR) and rule-based template matching dococr ; ha2022information , which are often ineffective for complex or varied layouts and fail to grasp the deep semantics of the content. LLMs, particularly Multimodal LLMs, have revolutionized document understanding by unifying the processing of text, layout, and visual information in three aspects: (1) Architectural innovations from the LayoutLM series to DocLLM, enable a synergistic understanding of document structure and content xu2020layoutlm ; xu2021layoutlmv2 ; huang2022layoutlmv3 ; wang2024docllm . Concurrently, (2) Retrieval-Augmented Generation (RAG) is employed for question answering and summarization across long or multiple documents lewis2020retrieval ; lei2024visdommultimodal , while (3) synthetic data generation effectively alleviates the scarcity of labeled data seol2024posterllama ; shi2025layoutcot . These techniques collectively drive the shift from simple information extraction to deep document reasoning and synthesis.
@@
$\bullet$ 3D Model. 3D models represent objects or scenes in Euclidean space, expressed as point clouds, meshes, or voxels, and are widely applied in scene understanding and scientific analysis. Conventional pipelines rely on geometric processing (e.g., mesh simplification, point cloud registration) hoppe1997view ; garland1997surface ; chen1992object and modeling software such as Blender or Maya blain2019complete ; tickoo2018autodesk , which require manual annotation and lack semantic understanding. Recent LLM-based approaches enable 3D–language alignment hong20233d , where geometry is transformed into structured embeddings or textual descriptions for reasoning. Agents orchestrate specialized 3D encoders and toolchains to support downstream tasks such as captioning, navigation, and scientific QA hong20233d ; xiong20253ur . Building on this foundation, 3D-Language Fusion provides unified frameworks that map point clouds and meshes into embeddings aligned with natural language, enabling tasks such as captioning and QA 3D-LLM hong20233d , 3UR-LLM xiong20253ur Domain-specific extensions integrate molecular and protein structures into multimodal reasoning, as in 3D-MoLM li2024towards , ProteinChat guo2023proteinchat , and ProtChatGPT wang2024protchatgpt . Moreover, 3D-Derived Task Enhancement leverages multi-agent systems leverage textual mediation of geometry for scene description, navigation, and retrieval, improving interpretability and efficiency hong20233d ; xiong20253ur ; li2024towards . Finally,Cross-Modal Refinement employs feature enhancement and domain adaptation techniques bridge 2D and 3D modalities, strengthening generalization in multimodal LLMs (e.g., visual grounding, 2D–3D alignment, or joint pretraining) hong20233d ; li2024towards ; li2024towards ; xiong20253ur .
@@
$\bullet$ From Tool-Coupled to Tool-Assisted. Traditional agents rely on tightly integrated, framework-specific tools. The evolution favors decoupled architectures, where LLMs can leverage arbitrary toolsets, enhancing flexibility and adaptability in diverse analytical contexts.
@@ 2 LLM for Structured Data Analysis
Takeaway: For relational data, LLM takes part in natural language interfaces, semantic analysis and time series analysis. For graph data, LLM can perform graph analysis query generation and semantic analysis.
@@ 2.1 Relational Data Analysis
LLM for Natural Language Interfaces. Basic analysis jobs for relational data are typically characterized by well-defined operations. These operations include basic calculations (e.g., summation, averaging, counting, ranking) relationaldata , statistical analysis (e.g., regression and clustering) hoff2015multilinear ; taskar2001probabilistic , and data quality assurance processes (e.g., constraint validation, outlier detection) maervoet2012outlier . Such tasks can generally be supported by tools like SQL or Python libraries (e.g., Pandas). Traditional methods typically rely on strictly defined SQL language and pre-defined analysis rules or algorithms. In contrast, LLMs can not only perform conventional analysis tasks, but also enable natural language–to–SQL/code conversion and advanced semantic analysis, leveraging strong language understanding and reasoning capabilities.
@@
(2) LLM-Based Analysis Agent: Data Interpreter hong2024datainterpreter leverages LLMs through APIs to generate task and action graphs. Specifically, they utilize LLM’s semantic reasoning ability to accurately decompose complex user queries into subproblems (e.g., correlation analysis, data exploration, and anomaly detection), and further refine and verify each subproblem to improve code generation results for data science tasks. Another approach bai2025collaboration fine-tunes a BART model DBLP:conf/acl/LewisLGGMLSZ20 as an intelligent agent with datasets constructed from task-instruction and role-definition prompts. The agent further generates more detailed prompts to guide LLMs to complete the NL2Code task step-by-step (e.g., role definition, demand optimization, code writing, and code review).
@@
LLM for Semantic Analysis. Moreover, some jobs require LLM-based analysis, such as those that involve semantic understanding or demand outputs in natural language format (e.g., table summarization). These challenges call for methodologies like (1) multi-step question answering (QA) with diverse decomposition strategies and (2) end-to-end QA leveraging specifically optimized LLMs.
@@
$\bullet$ Multi-Step QA. Multi-step question answering (QA) refers to decomposing complex queries into a sequence of sub-questions to facilitate step-by-step reasoning. According to the question decomposition mechanisms, existing methods can be categorized into two types: (1) static decomposition, which follows pre-defined and fixed processing steps (e.g., retrieve-select-reason), and (2) LLM-driven iterative decomposition, in which the LLM dynamically determines the next operation based on the contextual history of the reasoning process.
@@
(1) Static Decomposition: The static decomposition includes Retriever-Selector-Reasoner frameworks and the variants, which partition tasks into modular components for better multi-step inference and enhanced interpretability. The Extractor-Reasoner-Executor paradigm tatllm extracts the relevant segments from the context, generates the logic rules or equations, and performs the rules or executes the equations to get the final answer through LLM prompting. S3HQA lei2023s3hqa employs a retriever for initial filtering of heterogeneous resources, a selector for identifying the most relevant factual knowledge, and a generation-based reasoner for producing the final answer. For multi-table scenarios, SGAM wang2025plugging encodes schema links and join paths as a graph, transforms the query into interpretable reasoning chains on the graph, and executes along the chains to obtain the final answer.
(2) Iterative Decomposition: Static decomposition paradigm often struggles with multi-hop queries, while LLM-driven iterative decomposition overcomes these limitations by dynamically refining sub-tasks through recursive reasoning.
TAPERA zhao2024tapera introduces the query decomposition step into the question answering process by adopting the LLM-driven approach. A Planner first decomposes the query into sub-queries, forming an initial plan. The Reasoner then generates executable programs for each sub-query, and the Answer Generator derives answers based on the program outputs to fulfill the plan. Finally, the Planner updates or finalizes the plan as needed.
Similarly, ReAcTable zhang2023reactable and CHAIN-OF-TABLE wang2024chainoftable iteratively generate operations and update the table to present a reasoning chain as a proxy for intermediate thoughts through prompting LLMs and in-context learning.
@@
$\bullet$ End-to-End QA. End-to-End Question Answering (QA) refers to approaches in which the answer-generating LLM directly produces the final response without intermediate steps or iterative refinement. Based on the data representation and processing mechanisms, the relevant methods can be classified into table-specific LLM fine-tuning, table content retrieval, and table-as-image analysis.
@@
(1) Table-Specific LLM Fine-Tuning: Fine-tuning LLMs on task-specific table datasets enables them to internalize analytical knowledge directly within their parameters. TableGPT li2023tablegpt fine-tunes LLMs like GPT-3.5 using a diverse set of table tasks synthesized from real-world tables. Building on Qwen2.5 qwen2025qwen25 , TableGPT2 su2024tablegpt2 introduces a table encoder to generate a hybrid table representation, an adapter to generate query representations, and a LLM decoder to generate an agent workflow (i.e., the tool execution pipeline) for deriving the final answer. The TableGPT2 model is pre-trained on 593.8K tables and fine-tuned 2.36M question-answer pairs.
@@
(2) Table Content Retrieval: Instead of embedding the whole table, table content retrieval enhances model performance by eliminating noisy parts of the table while retaining information relevant to question answering. CABINET patnaik2024cabinet employs a weakly supervised component to produce a parsing statement that defines the criteria for selecting relevant rows and columns, emphasizing the corresponding table cell content. TableMaster cao2025tablemaster constructs a refined subtable through row and column lookup. By leveraging carefully designed LLM prompts (e.g., provide objective, table definition, table information, question, instructions, and response format), it ranks all candidate columns, selects a relevant subset based on the query, and then instructs the LLM to generate an SQL query for extracting the most relevant rows. In the test of Multi-table and Multi-hop Question Answering dataset (MMQA) wu2025mmqa , LLMs are required to retrieve tables from a given corpus and then perform multiple inferences. A multi-table retrieval method (MTR) that iteratively retrieves question-related and table-related tables achieves the best performance in the test.
@@
(3) Table-As-Image Analysis: Due to the limitations of (text-only) LLMs in understanding table structures, the table-as-image approach has been proposed, converting tables into images for analysis using multimodal LLMs. Table-LLaVA zheng2024multimodal applies incremental pretraining to LLaVA-7B liu2024llava on 150K table recognition samples (e.g., input a table image and output table representations in HTML, Markdown, or LaTeX), enabling the model to align table structures and elements with textual modality. It is further fine-tuned on 232K samples on question answering, text generation, fact verification, and structure understanding tasks to enhance its instruction-following ability. To enable a single model to perform various analytical tasks, TabPedia zhao2024tabpedia introduces the concept synergy mechanism, abstracting all table analysis tasks into concepts. Built on Vicuna-7B zheng2023judging , it appends meditative tokens to the input of the LLM decoder, which adaptively activates different regions of visual tokens and helps the model interpret the intent behind specific task questions. However, such methods face limitations when processing twisted or distorted tables, and their performance degrades significantly when directly handling document images.
LLM for Time Series Analysis. Time series data is a special type of relational data that can be stored in relational databases or in specialized time series databases like influxdb naqvi2017time . It represents the sequential data obtained by observing real or simulated scenes in chronological order, with each time step corresponding to one or more variables. Time series analysis typically aims to learn the trend or periodic pattern of data changes over time (e.g., sensor data, price fluctuation, traffic flow, etc.), and to model interdependence among variables (e.g., causality, correlation, coupling), enabling tasks such as prediction, classification, and detection liu2025towards . If viewed as a wide table indexed by time, time series analysis emphasizes the dynamic evolution of parameters along the timeline, in contrast to relational data analysis, which focuses on static relationships.
@@
Traditional time series analysis includes statistical-based and machine-learning-based methods, both rely on certain models such as ARIMA and LSTM 8614252 . However, statistical methods have shortcomings such as assuming linear relationships, tedious manual feature engineering, and limited ability for complex nonlinear patterns add rahmani2022association . The development of machine learning and deep learning provides a new solution for time series analysis 9005997 . At the same time, people are trying to improve the interpretability of analysis result 9713986 . Recently, LLMs for time series analysis have become a new focus due to their sequence modeling capabilities, less data dependencies, and the potential to capture long-term dependencies.
@@
$\bullet$ TS2NL. Due to the shared sequential features of time series data and natural language, the TS2NL method directly converts time series data into natural language and provides it to LLMs. The format of natural language brings better intuitiveness and interpretability.

### v2

@@ Abstract
Large language model (LLM) and agent techniques for data analysis (a.k.a LLM/Agent-as-Data-Analyst) have demonstrated substantial impact in both academica and industry. In comparison with traditional rule or small-model based approaches, (agentic) LLMs enable complex data understanding, natural language interfaces, semantic analysis functions, and autonomous pipeline orchestration. From a modality perspective, we review LLM-based techniques for $(i)$ structured data (e.g., table question answering for relational data and NL2GQL for graph data), $(ii)$ semi-structured data (e.g., markup languages understanding and semi-structured table modeling), $(iii)$ unstructured data (e.g., chart understanding, document understanding, code completion), and $(iv)$ heterogeneous data (e.g., data retrieval and modality alignment for data lakes). The technical evolution further distills five key design goals for intelligent data analysis agents, namely semantic-aware design, modality-hybrid integration, autonomous pipelines, tool-augmented workflows, and support for open-world tasks. Finally, we outline the remaining challenges and propose several insights and practical directions for advancing LLM/Agent-powered data analysis.
@@
Figure 2: Technical Overview of LLM/Agent-as-Data-Analyst. The five key design goals are illustrated in the center of the figure using distinct colors. The colored icons next to each technique indicate the specific design goal it supports.
@@ 1.2 Opportunities by LLM/Agent-As-Data-Analyst
O4: Autonomous Evolution. LLM agents function not as static tools but as adaptive systems capable of refining their performance through continuous interaction and learning. Such adaptability is essential, as analytical requirements and data environments inevitably evolve over time hong2024datainterpreter . For instance, a financial analysis agent may initially concentrate on trend forecasting but, through iterative feedback, expand its capabilities to detect anomalies in real time or incorporate emerging risk factors. This capacity for autonomous evolution ensures that the system remains advanced, alleviates the need for manual reconfiguration, and delivers increasingly personalized and efficient analytical support atreja2025alasautonomouslearningagent .
@@ 1.3 Techniques of LLM-Powered Data Analysis
As shown in Figure 1 and Figure 2, given the diversity of data formats and application scenarios, we establish a taxonomy of LLM-powered data analysis which progresses along two dimensions: $(i)$ the range of data modalities supported, i.e., structured, semi-structured, unstructured, and heterogeneous (the x-axis), and $(ii)$ the evolution of interaction paradigms, i.e., code-based, domain-specific language (DSL)-based, and natural language (NL)-based (the y-axis).
@@
LLM/Agent for Structured Data Analysis. Structured data, including relational databases relationaldata and graphs graphdata , remains central in industry for its standardized schema and well-defined semantics. Traditional methods rely on SQL or related query languages, later extended with DSLs for domain-specific tasks neo4j . With LLMs, users can interact with structured data via natural language through code generation, DSL mapping, or LLM-based question answering sqlsurvey ; shi2024sqlsurvey , while agent frameworks further orchestrate multi-step analytical workflows zhu2024sqlsurvey . The core of LLM-driven structured data analysis lies in employing LLMs for pipeline automation or end-to-end processing.
@@
$\bullet$ Relational Data. Relational data is typically stored and managed in specialized databases. An LLM-based approach to relational data analysis leverages the model as a natural language interface that translates analytical intents into SQL or executable code, which is then used to query the database. To enhance alignment between natural and formal languages, techniques such as schema linking zhang2024finsql , information retrieval DINSQL , and task decomposition hong2024datainterpreter are integrated into the pipeline, while task-specific fine-tuning yin2022code improves end-to-end generation. An alternative direction bypasses direct database manipulation, employing LLMs for semantic-level analysis. This requires deeper understanding of analytical intent and intra-table relationships. Accordingly, RAG patnaik2024cabinet , prompt engineering tatllm , and task decomposition zhao2024tapera are incorporated to enhance reasoning, while MLLMs zheng2024multimodal or LLMs su2024tablegpt2 are trained or fine-tuned on textual or visual table representations. Time-series data, a specialized subclass of relational data characterized by temporal dependencies, can similarly benefit from LLM-powered techniques, including natural language-to-code translation DBLP:conf/iclr/0005WMCZSCLLPW24 , sequence retrieval and transformation, and direct temporal reasoning alnegheimish2024can .
@@
$\bullet$ Graph Data. Graph data represents entities and their inter-dependencies to model complex network semantics, posing challenges due to its vast search space and intricate path reasoning. Specialized graph databases and query languages graphdb ; blazegraph enable direct access, while LLMs can serve as natural language interfaces for generating graph queries. Techniques such as agent-based frameworks liang2024natnl2gql , fine-tuning, and prompt engineering liang2024aligning enhance the model’s understanding of graph structures and query syntax. For semantic-level graph analysis, mainstream approaches leverage concepts from RAG zhang2022subgraph , agent-based reasoning luo2025kbqa , and fine-tuning ye2024instructglm to improve graph-aware comprehension and reasoning.
@@
$\bullet$ Semi-Structured Tables represent a flexible form of tabular data whose evolution is closely tied to advances in structured table understanding and the semantic capabilities of LLMs. Traditional approaches, often based on PLMs herzig2020tapas ; RoBERTa , struggle to capture the complex layouts, irregular headers, and hierarchical structures due to their limited ability to encode semantics. With the advent of LLMs, new paradigms have emerged for handling such tables, including semi-structured to structured transformations that convert semi-structured tables into relational-like forms 10.1145/3726302.3730071 ; li2023autotablessynthesizingmultisteptransformations , and DSL-based modeling that leverages domain-specific languages to explicitly encode structure and operations tang2025st .
LLM/Agent for Unstructured Data Analysis. Unstructured data encompass a wide range of modalities such as charts, videos, documents, and 3D models, which lack a fixed schema and therefore pose significant challenges for conventional analysis pipelines.
@@
$\bullet$ Video. Video encodes spatial content that evolves over time, requiring joint modeling of semantics and dynamics. Conventional methods rely on vision backbones with temporal pooling or attention, which struggle with annotation costs and long-sequence efficiency chen2024timemarker wang2024grounded . Recent advances reformulate videos as structured token sequences for LLM reasoning. Agents further decompose queries into temporal grounding, multimodal fusion, and summarization, enabling richer temporal understanding and efficient computation chen2024timemarker deng2025seq2time . Building upon this, LLMs have also been extended to Video Emotional Analysis , where multimodal fusion of visual, acoustic, and textual cues enables affective state inference.In addition, posture-based emotion analysis and 3D mesh reconstruction support nuanced social interaction modeling and relational emotional predictions jadhav2023ai ; muller2024predicting . For Object Detection, integrating high-precision detectors with multimodal reasoning allows object-centric summarization and reference grounding in videos to make correct object detections in videos de2024video ; yuan2025videorefer . Furthermore, Gesture and Behavior Detection leverages LLM-driven pipelines to extract fine-grained motion features and gestures, supporting interaction analysis and embodied behavior reasoning, often via 3D reconstructions for higher fidelity whitehead2025utilizing ; weng2025artificial .
$\bullet$ Document. Documents, such as PDFs, web pages, and scanned reports, serve as primary carriers of information in business and academia. Traditional document analysis relies on Optical Character Recognition (OCR) and rule-based template matching dococr ; ha2022information , which are often ineffective for complex or varied layouts and fail to grasp the deep semantics of the content. LLMs, particularly Multimodal LLMs, have revolutionized document understanding by unifying the processing of text, layout, and visual information in three aspects: $(i)$ Architectural innovations from the LayoutLM series to DocLLM, enable a synergistic understanding of document structure and content xu2020layoutlm ; xu2021layoutlmv2 ; huang2022layoutlmv3 ; wang2024docllm . Concurrently, $(ii)$ Retrieval-Augmented Generation (RAG) is employed for question answering and summarization across long or multiple documents lewis2020retrieval ; lei2024visdommultimodal , and $(iii)$ synthetic data generation effectively alleviates the scarcity of labeled data seol2024posterllama ; shi2025layoutcot . These techniques collectively drive the shift from simple information extraction to deep document reasoning and synthesis.
@@
$\bullet$ 3D Model. 3D models represent objects or scenes in Euclidean space, expressed as point clouds, meshes, or voxels, and are widely applied in scene understanding and scientific analysis. Conventional pipelines rely on geometric processing (e.g., mesh simplification, point cloud registration) hoppe1997view ; garland1997surface ; chen1992object and modeling software such as Blender or Maya blain2019complete ; tickoo2018autodesk , which require manual annotation and lack semantic understanding. Recent LLM-based approaches enable 3D–language alignment hong20233d , where geometry is transformed into structured embeddings or textual descriptions for reasoning. Agents orchestrate specialized 3D encoders and toolchains to support downstream tasks such as captioning, navigation, and scientific QA hong20233d ; xiong20253ur . Building on this foundation, 3D-Language Fusion provides unified frameworks that map point clouds and meshes into embeddings aligned with natural language, enabling tasks such as captioning and QA 3D-LLM hong20233d , 3UR-LLM xiong20253ur Domain-specific extensions integrate molecular and protein structures into multimodal reasoning, as in 3D-MoLM li2024towards , ProteinChat guo2023proteinchat , and ProtChatGPT wang2024protchatgpt . Moreover, 3D-Derived Task Enhancement leverages multi-agent systems leverage textual mediation of geometry for scene description, navigation, and retrieval, improving interpretability and efficiency hong20233d ; xiong20253ur ; li2024towards . Finally, Cross-Modal Refinement employs feature enhancement and domain adaptation techniques(e.g., visual grounding, 2D–3D alignment, or joint pretraining) to bridge 2D and 3D modalities, strengthening generalization in multimodal LLMs hong20233d ; li2024towards ; li2024towards ; xiong20253ur .
@@
$\bullet$ From Tightly-Coupled to Flexible Tools. Traditional agents rely on tightly integrated, framework-specific tools. The evolution favors decoupled architectures, where LLMs can leverage arbitrary toolsets, enhancing flexibility and adaptability in diverse analytical contexts.
@@ 2.1 Relational Data Analysis
LLM for Natural Language Interfaces. Relational data analysis typically involves well-defined operations such as aggregation (e.g., summation, averaging, ranking) relationaldata , statistical modeling (e.g., regression, clustering) hoff2015multilinear ; taskar2001probabilistic , and data quality assurance (e.g., constraint validation, outlier detection) maervoet2012outlier , often supported by SQL or Python libraries like Pandas. While traditional methods depend on rigid SQL syntax and predefined analytical rules, LLMs extend these capabilities by enabling natural language to SQL/code translation and advanced semantic reasoning through their strong language understanding and inference abilities.
@@
(2) LLM-Based Analysis Agent: Data Interpreter hong2024datainterpreter employs LLMs via APIs to generate task and action graphs, leveraging their semantic reasoning to decompose complex user queries (e.g., correlation analysis, data exploration, anomaly detection) into subproblems and iteratively refine and verify each to improve code generation for data science tasks. Another approach bai2025collaboration fine-tunes a BART model DBLP:conf/acl/LewisLGGMLSZ20 as an intelligent agent using datasets derived from task-instruction and role-definition prompts. The agent produces detailed prompts to guide LLMs through stepwise NL2Code execution, including role definition, requirement optimization, code writing, and code review.
@@
LLM for Semantic Analysis. Moreover, tasks requiring semantic understanding or natural language outputs (e.g., table summarization) benefit from LLM-based methods, including $(i)$ multi-step question answering with decomposition strategies and $(ii)$ end-to-end QA using optimized LLMs.
@@
$\bullet$ Multi-Step QA. Multi-step question answering decomposes complex queries into sequential sub-questions for stepwise reasoning. Existing methods include $(i)$ static decomposition, following fixed processing steps (e.g., retrieve–select–reason), and $(ii)$ LLM-driven iterative decomposition, where the LLM dynamically determines each step based on contextual reasoning history.
@@
(1) Static Decomposition: Static decomposition frameworks, such as Retriever–Selector–Reasoner and its variants, divide tasks into modular components to support multi-step inference and improve interpretability. The Extractor–Reasoner–Executor paradigm tatllm extracts relevant context, generates logic rules or equations, and executes them via LLM prompting to get the final answer. Similarly, S3HQA lei2023s3hqa uses a retriever to filter heterogeneous resources, a selector to identify relevant knowledge, and a generation-based reasoner to produce the final answer. For multi-table scenarios, SGAM wang2025plugging encodes schema links and join paths as a graph, converts queries into reasoning chains on the graph, and executes them to produce the final answer.
(2) Iterative Decomposition: Static decomposition often struggles with multi-hop queries, whereas LLM-driven iterative decomposition dynamically overcomes it by refining sub-tasks via recursive reasoning. TAPERA zhao2024tapera applies this approach by introducing query decomposition, where the Planner breaks the query into sub-queries, the Reasoner generates executable programs for each, and the Answer Generator derives results to fulfill the plan. Finally, the Planner updates or finalizes the plan as needed. Similarly, ReAcTable zhang2023reactable and CHAIN-OF-TABLE wang2024chainoftable iteratively generate operations and update tables, using LLMs and in-context learning to produce reasoning chains as proxies for intermediate thoughts.
@@
$\bullet$ End-to-End QA. End-to-End Question Answering refers to methods where the LLM directly generates the final answer without intermediate steps or iterative refinement. Approaches are classified by data handling into table-specific LLM fine-tuning, table content retrieval, and table-as-image analysis.
@@
(1) Table-Specific LLM Fine-Tuning: Fine-tuning LLMs on task-specific table datasets embeds analytical knowledge directly into model parameters. TableGPT li2023tablegpt fine-tunes GPT-3.5 on diverse real-world table tasks, while TableGPT2 su2024tablegpt2 , built on Qwen2.5 qwen2025qwen25 and pre-trained on 593.8K tables and fine-tuned on 2.36M QA pairs, introduces a table encoder for hybrid table representation, an adapter for query encoding, and an LLM decoder to generate agent workflows like tool execution pipelines for deriving the final answer.
@@
(2) Table Content Retrieval: Rather than embedding entire tables, content retrieval selects relevant rows and columns to reduce noise. CABINET patnaik2024cabinet employs a weakly supervised component to produce a parsing statement that defines the criteria for selecting relevant cells, while TableMaster cao2025tablemaster refines subtables via row and column lookup, ranks candidate columns using LLM prompts, selects query-relevant subsets, and generates SQL queries to extract the most pertinent rows. On the Multi-table and Multi-hop QA dataset (MMQA) wu2025mmqa , LLMs must retrieve relevant tables from a corpus and perform multiple inferences. Iterative multi-table retrieval (MTR), fetching question- and table-related tables stepwise, achieves the best performance.
@@
(3) Table-As-Image Analysis: To overcome text-only LLM limitations in understanding table structures, tables are converted to images for multimodal LLM analysis. Table-LLaVA zheng2024multimodal pretrains LLaVA-7B liu2024llava on 150K table recognition samples (e.g., output HTML of the input table) to align table structures and elements with textual modality and fine-tunes on 232K downstream tasks (e.g., fact verification, question answering) to enhance its instruction-following ability. TabPedia zhao2024tabpedia enables a single model to handle diverse table analysis tasks via a concept-synergy mechanism that abstracts tasks into concepts. Built on Vicuna-7B zheng2023judging , it adds meditative tokens to the LLM decoder input, adaptively activating visual token regions to interpret task-specific intents. However, limitations remain for such methods when processing twisted, distorted tables and direct document-image inputs.
LLM for Time Series Analysis. Time series data, a specialized form of relational data stored in relational or time-series databases (e.g., InfluxDB naqvi2017time ), captures sequential observations over time, with each step recording one or more variables. Analysis focuses on temporal trends, periodic patterns, and inter-variable dependencies (e.g., causality, correlation, coupling) to support prediction, classification, and detection liu2025towards . Unlike static relational analysis, time series emphasizes the dynamic evolution of variables along the timeline.
@@
Traditional time series analysis relies on statistical and machine-learning methods, such as ARIMA and LSTM 8614252 . However, statistical approaches are limited by linear assumptions, manual feature engineering, and difficulty capturing complex nonlinear patterns rahmani2022association . Advances in machine and deep learning offer more flexible solutions 9005997 , alongside efforts to improve result interpretability 9713986 . Recently, LLMs have emerged for time series analysis, leveraging their sequence modeling, generalization capacity, and ability to capture long-term dependencies.
@@
$\bullet$ TS2NL. Due to the sequential similarity between time series and natural language, TS2NL converts time series data into natural language for LLMs, enhancing intuitiveness and interpretability.